OpenStudy (anonymous):

How to solve |3x-1|-4≥3 I got x=8/3 ... I tried plugin it in, but it didnt work

OpenStudy (anonymous):

\[\left| 3x-1 \right|\geq 3+4\]

OpenStudy (anonymous):

\[\left| 3x-1 \right| \geq 7,3x-1\leq -7 and ~3x-1 \geq7\] find x for both.